# Speech-Language Models with Decoupled Tokenizers and Multi-Token Prediction

### v1

@@ 4.5 Ablation Study
MTP’s Architecture Selection We analyze fusion strategies and speech head architectures to select the optimal MTP configuration, as summarized in Table 6. NTP’s Layer and Token Organization Selection We explore the best number of decoupled tokenizer codebook layers and speech token organization patterns for SLM training. Details can be found in Section E.2 and Table 8.
@@ 4.6 Effect of MTP on Cross-modal Alignment
We include complementary qualitative results in Section F.1.
@@ 4.7 Effect of Tokenizer-MTP Combination
To examine the effectiveness of different tokenizer coupling strategies in MTP SLMs, we perform comprehensive experiments in Section F.2.
@@ 5 Related Work
have advanced through RVQ-based architectures and quantization innovations: SoundStream (2022) introduced the RVQ-VAE framework, enhanced by Encodec (2023, LSTM temporal modeling) and SpeechTokenizer (2024, HuBERT semantic distillation). Recent ultra-low-bitrate methods include Improved RVQGAN (DAC, 2023) with quantizer dropout, WavTokenizer (2025, 0.9kbps via single-layer RVQ), BigCodec (2024, 1.04kbps via Convolution-LSTM projection), and StableCodec (2024, 0.7kbps using Transformer with FSQ (2024)).
@@
LLMs are increasingly adopted for unified speech processing through two primary paradigms. The first integrates speech encoders with LLMs for comprehension tasks like ASR Ma et al. (2024) and spoken language understanding Chu et al. (2023), achieving strong analytical capabilities but lacking generative functionality. The second targets speech generation via dual strategies: codec-based models (e.g., SpeechGPT, 2023) discretize speech into compressed tokens, while diffusion-enhanced approaches (e.g., CosyVoice, 2024a; 2024c) synthesize waveforms through LLM-diffusion.
@@
has emerged as a critical technique in LLMs to accelerate inference. Detailed related works are placed on Appendix H.
@@ 6 Conclusion
Limitations
Verification on Other Decoupled Speech Tokenizers
While our decoupled SLM framework shows promising results, broader validation across comparable systems remains advisable. The current evaluation is limited to FACodec as no more open-source public implementations can be found in this emerging field to the best of our knowledge. Future work can extend validation to other decoupled speech tokenizers to better assess generalizability of our method.
Challenges under High Compression
Although MTP consistently improves cross-modal alignment and reduces the Word Error Rate (WER), achieving near ground truth, even at 12× compression. However, this aggressive compression results in a slight decline in UTMOS. This quality degradation probably stems from the loss of fine-grained acoustic details at a high compression ratio. Our future work will investigate adaptive compression strategies or richer intermediate representations to better enhance the synthesis quality.
@@ A.1 Compared Baselines
Encodec Défossez et al. (2023): An RVQ-based codec designed for universal audio compression.666https://huggingface.co/facebook/encodec_24khz
WavTokenizer Ji et al. (2025): A single VQ codebook-based tokenizer for universal audio.777https://huggingface.co/novateur/WavTokenizer-large-speech-75token
StableCodec Parker et al. (2024): A residual FSQ-based tokenizer for speech.888https://huggingface.co/stabilityai/stable-codec-speech-16k
BigCodec Xin et al. (2024): A VQ-based single-stream codec for speech.999https://huggingface.co/Alethia/BigCodec/resolve/main/bigcodec.pt
SpeechTokenizer Zhang et al. (2024): An RVQ-based codec with semantic distillation for speech.101010https://huggingface.co/fnlp/AnyGPT-speech-modules/tree/main/speechtokenizer
@@ A.2 Results of Codec Reconstruction
Model |
Nq |
STOI |
UTMOS | SIM | |||||||
Ground Truth | - | - | - | - | 1.00 | 4.55 | 4.09 | 1.00 | |||||||
Encodec | 1024 | 8 | 600 | 6000 | 0.94 | 3.18 | 3.09 | 0.88 | |||||||
WavTokenizer | 4096 | 1 | 75 | 900 | 0.90 | 2.63 | 3.79 | 0.65 | |||||||
StableCodec | 15625 | 2 | 50 | 697 | 0.91 | 2.91 | 4.23 | 0.62 | |||||||
BigCodec | 8192 | 1 | 80 | 1040 | 0.93 | 3.27 | 4.11 | 0.84 | |||||||
SpeechTokenizer | 1024 | 8 | 100 | 1000 | 0.92 | 2.61 | 3.87 | 0.83 | |||||||
FACodec | 1024 | 6 | 480 | 4800 | 0.95 | 3.47 | 3.90 | 0.87 |
Table 5 shows the results of VQ reconstruction evaluation on the LibriTTS test-clean dataset. The results reveal that for acoustic quality (STOI Taal et al. (2010), PESQ Rix et al. (2001) and UTMOS), all tokenizers perform similar with the ground truth, sharing close metric values. But for SIM, the trend differs: the VQ performs better on SIM as its token rate increases, and the decoupled model (FACodec) surpasses most of the coupled and semi-decoupled models. This is probably because reduced token rates compromise timbre fidelity, combined with coupled modeling diminishing acoustic prominence relative to semantic components.
@@ Appendix C Experimental Setups
We selected the current state-of-the-art speech tokenizer as the baseline for our model. To ensure a fair comparison, we employed the official weight files provided by the Encodec Défossez et al. (2023), SpeechTokenizer Zhang et al. (2024), WavTokenizer Ji et al. (2025), StableCodec Parker et al. (2024), BigCodec Xin et al. (2024). Details can be found in Section A.1.
@@
Our baseline training configuration employs the AdamW optimizer ( and ) with a 5e-4 initial learning rate following a cosine decay schedule. All models undergo 90k training steps on 8 A800 GPUs using consistent batch sizes of 16 for fair comparison. For specialized tasks like Role-Playing Knowledge QA, we implement a two-phase adaptation: 40k steps for cross-modal alignment pretraining followed by 12k steps of knowledge-role joint fine-tuning. During inference, we apply greedy decoding with a 1.2 repetition penalty to balance output quality and diversity.
@@
For the default TTS evaluation, following Zhang et al. (2024), we randomly sample 300 samples from the LibriTTS test set for speech synthesis evaluation. For Role-Playing Knowledge QA, the evaluation set of RoleTriviaQA contains 1,500 in-domain and 926 out-of-domain samples.
@@
For the default TTS evaluation, we evaluate Success Rate (SR), speech quality (UTMOS Saeki et al. (2022)), Word Error Rate (WER), and speaker similarity (SIM).
For Role-Playing Knowledge QA task, we report three metrics: Exact Match (EM) and F1 score for text responses, while SIM for speech responses. Details of these metrics can be found in Appendix D.
@@ Appendix D Details of Evaluation Metrics
UTMOS assesses the acoustic quality of the generated speech. We utilized the model provided by SaruLab121212https://huggingface.co/spaces/sarulab-speech/UTMOS-demo/resolve/main/epoch%3D3-step%3D7459.ckpt to evaluate it.
Speaker Similarity (SIM) reflects how similar it is between the generated speech and the reference speech. We considered the WavLM-TDCNN Chen et al. (2022) speaker embedding model to assess speaker similarity.
2 more metrics are considered for the evaluation of the VQ reconstruction (See Appendix A.2):
@@
Appendix E Ablation Study
E.1 Effect of MTP Architecture
@@
E.2 Effect of Number of Codebook Layers in SLMs
@@
E.3 Effect of Speech Token Organization Pattern
@@
Table 8 shows that FWI outperforms CWI on all metrics especially on WER (6.07 vs. 22.21), which means the CWI-trained model aligns text and speech modalities less effectively than its FWI-trained counterpart. Therefore, we opt FWI as the default setting for our main experiments.
Appendix F Analysis
F.1 Quantitative Analysis of MTP’s Impact on Cross-Modal Alignment
@@
F.2 Effect of Tokenizer-MTP Combination
@@
Appendix G Details of RoleTriviaQA
@@ G.1 Data Construction
Table 10 demonstrates the detailed composition of RoleTriviaQA along with the selected speakers, with the ID evaluation samples are collected from TriviaQA test set and OOD samples are gathered from Web Questions of OpenAudioBench Li et al. (2025). Notably, we added about 26k text QA training samples filtered from LLaVA-Instruct-150k131313https://huggingface.co/datasets/liuhaotian/LLaVA-Instruct-150K during training to preserve the original textual ability of the SLM.
@@
For each character’s speech synthesis, we select high-quality reference speeches by combining speaker embedding similarity metrics with manual filtering as an comprehensive criteria. Then we leverage the CosyVoice 2 141414https://huggingface.co/FunAudioLLM/CosyVoice2-0.5B system to generate role-playing speech from text answers. Finally, all speeches are resampled to 16 kHz to fit the tokenization requirement.
G.2 Samples of RoleTriviaQA
G.3 Prompt template used in Knowledge-Role Joint Fine-Tuning
Appendix H Multi Token Prediction
Multi Token Prediction (MTP) is leveraged in LLMs to accelerate inference through three principal strategies: 1) Multi-head architectures enhance generation quality via parallel output heads Gloeckle et al. (2024); 2) Tree-based systems like MEDUSA Cai et al. (2024) refine candidate selection through hierarchical validation; 3) Scalable frameworks such as DeepSeek-V3 DeepSeek-AI et al. (2024) enable efficient large-scale training. Recent adaptations extend MTP to speech processing: SLAM-Omni Chen et al. (2024a) employs audio token grouping for computational efficiency but neglects paralinguistic modeling through semantic-focused processing, while VocalNet Wang et al. (2025) achieves localized acceleration via DeepSeek-V3-style parallel prediction yet faces pipeline complexity from flow-matching requirements.

### v2

@@ 4.5 Ablation Study
MTP Architecture Selection
We analyze 6 different fusion strategies and speech head architectures to select the optimal MTP configuration (MLP-Linear combination), as elaborated in Table 8.
NTP Layer and Token Organization Selection
We explore the best number of decoupled tokenizer codebook layers and speech token organization patterns for SLM training. Results show that 3 codebook layers and Frame-Wise Interleaving organization pattern is the best-performing setting. Details are provided in Section F.2 and Section F.3.
RoleTriviaQA Data Format Selection
@@ 4.6 Effect of MTP on Cross-modal Alignment
We include complementary qualitative results in Section G.1.
@@ 4.7 Effect of Tokenizer-MTP Combination
To examine the effectiveness of different tokenizer coupling strategies in MTP SLMs, we perform comprehensive experiments in Section G.2.
@@ 5 Related Work
has advanced through RVQ-based architectures and quantization innovations: SoundStream (2022) introduced the RVQ-VAE framework, enhanced by Encodec (2023, LSTM temporal modeling) and SpeechTokenizer (2024, HuBERT semantic distillation). Recent ultra-low-bitrate methods include Improved RVQGAN (DAC, 2023) with quantizer dropout, WavTokenizer (2025, 0.9kbps via single-layer RVQ), BigCodec (2024, 1.04kbps via Convolution-LSTM projection), and StableCodec (2024, 0.7kbps using Transformer with FSQ (2024)).
@@
LLMs are increasingly adopted for unified speech processing through two primary paradigms. The first integrates speech encoders with LLMs for comprehension tasks like ASR (Ma et al. 2024) and spoken language understanding (Chu et al. 2023), achieving strong analytical capabilities but lacking generative functionality. The second targets speech generation via dual strategies: codec-based models (e.g., SpeechGPT, 2023) discretize speech into compressed tokens, while diffusion-enhanced approaches (e.g., CosyVoice, 2024a; 2024c) synthesize waveforms through LLM-diffusion.
@@
has emerged as a critical technique in LLMs to accelerate inference. Detailed related works are placed on Appendix I.
@@ A.1 Compared Baselines
Encodec (Défossez et al. 2023): An RVQ-based codec designed for universal audio compression.666https://huggingface.co/facebook/encodec˙24khz
WavTokenizer (Ji et al. 2025): A single VQ codebook-based tokenizer for universal audio.777https://huggingface.co/novateur/WavTokenizer-large-speech-75token
StableCodec (Parker et al. 2024): A residual FSQ-based tokenizer for speech.888https://huggingface.co/stabilityai/stable-codec-speech-16k
BigCodec (Xin et al. 2024): A VQ-based single-stream codec for speech.999https://huggingface.co/Alethia/BigCodec/resolve/main/bigcodec.pt
SpeechTokenizer (Zhang et al. 2024): An RVQ-based codec with semantic distillation for speech.101010https://huggingface.co/fnlp/AnyGPT-speech-modules/tree/main/speechtokenizer
@@ A.2 Results of Codec Reconstruction
Table 5 shows the results of VQ reconstruction evaluation on the LibriTTS test-clean dataset. The results reveal that for acoustic quality (STOI (Taal et al. 2010), PESQ (Rix et al. 2001) and UTMOS), all tokenizers perform similar with the ground truth, sharing close metric values. But for SIM, the trend differs: the VQ performs better on SIM as its token rate increases, and the decoupled model (FACodec) surpasses most of the coupled and semi-decoupled models. This is probably because reduced token rates compromise timbre fidelity, combined with coupled modeling diminishing acoustic prominence relative to semantic components.
@@ Appendix C Experimental Setups
We selected the current state-of-the-art speech tokenizer as the baseline for our model. To ensure a fair comparison, we employed the official weight files provided by the Encodec (Défossez et al. 2023), SpeechTokenizer (Zhang et al. 2024), WavTokenizer (Ji et al. 2025), StableCodec (Parker et al. 2024), BigCodec (Xin et al. 2024). Details can be found in Section A.1.
@@
Our baseline training configuration employs the AdamW optimizer ( and ) with a 5e-4 initial learning rate following a cosine decay schedule. All models undergo 90k training steps using consistent batch sizes of 16 for fair comparison. For specialized tasks like Role-Playing Knowledge QA, we implement a two-phase adaptation: 40k steps for cross-modal alignment pretraining followed by 12k steps of knowledge-role joint fine-tuning. During inference, we apply greedy decoding with a 1.2 repetition penalty to balance output quality and diversity.
Hardware and Software Requirements
All experiments are conducted with 8*NVIDIA A800 GPUs and CUDA version 12.4 on Ubuntu 20.04.6 LTS.
@@
For the default TTS evaluation, following (Zhang et al. 2024), we randomly sample 300 samples from the LibriTTS test set for speech synthesis evaluation. For Role-Playing Knowledge QA, the evaluation set of RoleTriviaQA contains 1,500 in-domain and 926 out-of-domain samples.
@@
For the default TTS evaluation, we evaluate Success Rate (SR), speech quality (UTMOS (Saeki et al. 2022)), Word Error Rate (WER), and speaker similarity (SIM).
For Role-Playing Knowledge QA task, we report three metrics: Exact Match (EM) and F1 score for text responses, while SIM for speech responses. Details of these metrics can be found in Table 6.
@@ Appendix D Details of Evaluation Metrics
Model | Model Size | Data Format | EM | F1 |
SpiritLM(2024) | 7B | SQ, SA | 0.0 | 2.0 |
Moshi(2024) | 7.6B | SQ, SA | 0.0 | 2.8 |
Mini-Omni(2024) | 0.5B | TQ, SA | 0.0 | 8.2 |
SpiritLM | 7B | TQ, SA | 0.0 | 3.9 |
GLM-4-Voice(2024) | 9B | TQ, SA | 0.1 | 14.9 |
SpeechGPT(2023) | 7B | SQ, TQ, TA, SA | 0.0 | 6.6 |
Best Baseline | 0.5B | TQ, TA, SA | 12.0 | 23.8 |
@@
UTMOS assesses the acoustic quality of the generated speech. We utilized the model provided by SaruLab121212https://huggingface.co/spaces/sarulab-speech/UTMOS-demo/tree/main to evaluate it.
Speaker Similarity (SIM) reflects how similar it is between the generated speech and the reference speech. We considered the WavLM-TDCNN (Chen et al. 2022) speaker embedding model to assess speaker similarity.
2 more metrics are considered for the evaluation of the VQ reconstruction (See Appendix 4):
@@
Appendix E RoleTriviaQA Results on SLM Baselines
@@
Results in Table 6 show that the best baseline (SLM trained with decoupled tokenizer FACodec) in our study reaches remarkably higher EM and F1 (12.0 and 23.8) than other SLM baselines, most of which fail to generate exactly correct answers (EM=0). This result confirms the superiority of decoupled SLMs to other baselines with different architechtures.
Appendix F Ablation Study
F.1 Effect of MTP Architecture
@@
F.2 Effect of Number of Codebook Layers in SLMs
@@ F.3 Effect of Speech Token Organization Pattern
Table 9 shows that FWI outperforms CWI on all metrics especially on WER (6.07 vs. 22.21), which means the CWI‐trained model aligns text and speech modalities less effectively than its FWI‐trained counterpart. Therefore, we opt FWI as the default setting for our main experiments.
F.4 Effect of RoleTriviaQA data format
Data Format | Model | ID | OOD | ||||
EM | F1 | SIM | EM | F1 | SIM | ||
{SQ, SA} | StableCodec-NTP-1H | 0.0* | 0.6* | 0.21 | 0.1* | 0.2* | 0.21 |
FACodec-NTP-1H | 0.5* | 2.4* | 0.60 | 0.4* | 1.0* | 0.60 | |
FACodec-MTP-12H | 0.6* | 2.5* | 0.60 | 0.3* | 1.6* | 0.59 | |
{TQ, SA} | StableCodec-NTP-1H | 0.0* | 0.3* | 0.17 | 0.0* | 0.2* | 0.16 |
FACodec-NTP-1H | 0.5* | 2.4* | 0.60 | 0.8* | 1.6* | 0.60 | |
FACodec-MTP-12H | 0.7* | 3.1* | 0.59 | 0.3* | 1.6* | 0.59 | |
{TQ, TA, SA} | StableCodec-NTP-1H | 7.5 | 13.8 | 0.17 | 0.6 | 3.6 | 0.16 |
FACodec-NTP-1H | 12.0 | 23.8 | 0.61 | 9.0 | 17.9 | 0.61 | |
FACodec-MTP-12H | 18.0 | 29.6 | 0.61 | 10.8 | 21.4 | 0.60 |
We evaluate three data formats for RoleTriviaQA:
(i.e., {SQ, SA})
(i.e., {TQ, SA})
(i.e., {TQ, TA, SA})
Table 10 compares these formats for SLMs trained with StableCodec and FACodec under both NTP and MTP settings131313Speaker embedding is unavailable for coupled tokenizers..
Results show that the TQ, TA, SA format performs best, achieving top scores (EM 18.0, F1 29.6) with FACodec and MTP heads. It also significantly improves performance for StableCodec-trained SLMs compared to other formats (EM 7.5 vs. 0.0; F1 13.8 vs. 0.6).
Appendix G Analysis
G.1 Quantitative Analysis of MTP’s Impact on Cross-Modal Alignment
@@
G.2 Effect of Tokenizer-MTP Combination
@@
Appendix H Details of RoleTriviaQA
@@ H.1 Data Construction
Table 12 demonstrates the detailed composition of RoleTriviaQA along with the selected speakers. The ID evaluation samples are collected from TriviaQA test set and OOD samples are gathered from Web Questions subset of OpenAudioBench. Notably, we added about 26,000 text QA training samples filtered from LLaVA-Instruct-150k141414https://huggingface.co/datasets/liuhaotian/LLaVA-Instruct-150K during training to preserve the original textual ability of the SLM.
@@
For each character’s speech synthesis, we select high-quality reference speeches by combining speaker embedding similarity metrics with manual filtering as an comprehensive criteria. Then we leverage the CosyVoice 2 151515https://huggingface.co/FunAudioLLM/CosyVoice2-0.5B system to generate role-playing speech from text answers. Finally, all speeches are resampled to 16 kHz to fit the tokenization requirement.
H.2 Samples of RoleTriviaQA
H.3 Prompt template used in Knowledge-Role Joint Fine-Tuning
Appendix I Multi Token Prediction
Multi Token Prediction (MTP) is leveraged in LLMs to accelerate inference through three principal strategies: 1) Multi-head architectures enhance generation quality via parallel output heads (Gloeckle et al. 2024); 2) Tree-based systems like MEDUSA (Cai et al. 2024) refine candidate selection through hierarchical validation; 3) Scalable frameworks such as DeepSeek-V3 (DeepSeek-AI et al. 2024) enable efficient large-scale training. Recent adaptations extend MTP to speech processing: SLAM-Omni (Chen et al. 2024a) employs audio token grouping for computational efficiency but neglects paralinguistic modeling through semantic-focused processing, while VocalNet (Wang et al. 2025) achieves localized acceleration via DeepSeek-V3-style parallel prediction yet faces pipeline complexity from flow-matching requirements.